# Analyzing Multifiltering Functions Using Multiparameter Discrete Morse Theory

### v1

@@ 3.3. Critical components
Let $R$ be a relation on the set ${\mathcal{C}}$ of critical simplices of a mdm function $g:K\rightarrow{\mathbb{R}}^{k}$ and consider its transitive closure $\bar{R}$. Namely, $\sigma\bar{R}\tau$ if there is a sequence $\sigma=\sigma_{0},\sigma_{1},...,\sigma_{n}=\tau\in{\mathcal{C}}$ such that $\sigma_{i-1}R\sigma_{i}$ for each $i=1,...,n$. If $R$ is reflexive and symmetric, it is easy to prove that $\bar{R}$ is an equivalence relation $\sim_{R}:=\bar{R}$, so it can be used to partition ${\mathcal{C}}$. We then call a critical component of $g$ (with respect to $\sim_{R}$) an equivalence class of $\sim_{R}$.
As an example, if $R$ is such that $\sigma R\tau\Leftrightarrow\sigma\leq\tau\text{ or }\sigma\geq\tau$, then $\sim_{R}$ yields the partition of ${\mathcal{C}}$ into connected components (with respect to the Alexandrov topology). As mentioned above, connectedness is too strict of a criteria to define critical components. Thus, we instead consider the following definition, which was proposed in [14]. Namely, let $\sim_{g}:=\bar{R}_{g}$, where $R_{g}$ is such that $\sigma R_{g}\tau$ if
@@ 2. $\mathrm{(2)}$
The formal definition of $\rightarrowtriangle_{g}$ is not presented here, but the idea is that $\sigma\rightarrowtriangle_{g}\tau$ when $\sigma$ is connected to $\tau$ in the flow induced by the gradient field ${\mathcal{V}}$ of $g$. This is the case, for example, when $\sigma\geq\tau$ or when there is a ${\mathcal{V}}$-path going from a face of $\sigma$ to a coface of $\tau$.
@@
For toy examples, defining critical components using $\sim_{g}$ yields desirable results, as argued in [14]. For instance, the mdm function from Example 3.3 has two critical components with respect to $\sim_{g}$, represented in red and orange in Figure 3. In Section 7.3, we will see how this definition of critical components of a mdm function performs on a larger scale.
@@ 4.1. Description
First of all, apart from the complex $K$ and the map $f$, the main algorithm takes two input variables: an injective real-valued map $I:K\rightarrow{\mathbb{R}}$, called an indexing map, and a small parameter $\epsilon>0$. The indexing map $I$ is used to order the simplices of $K$ while the parameter $\epsilon$ is used to ensure the mdm function $g$ computed is such that, for all $\sigma\in K$ and all $i=1,...,k$,
@@
$w_{1}$ $=\max\left(\{f_{1}(\sigma)\}\bigcup_{\alpha\in A}\{g_{1}(\alpha)\}\right),$
@@
7:     $w_{1}=\max\{f_{1}(\sigma)\}\bigcup_{\alpha\in A}\{g_{1}(\alpha)\}$
@@ 4.2. Complexity analysis
  • The index mapping $I$ is precomputed. As explained previously, the computation of $I$ can be included in the implementation of $K$.
@@
Moreover, the complexity of GenerateMDM depends greatly on the execution of ExpandMDM. For each level set $L_{u}$, we will see in Section 5 that each $\sigma\in L_{u}$ is added to each priority queue PQzero and PQone at most once and is processed exactly once. We compute the cost of processing each $\sigma\in L_{u}$ by considering each step of its process separately.
@@ 3. $\mathrm{(3)}$
    From Algorithm 3, we see that to define $g(\sigma)$, in the worst case, ComputeG needs to compute $\max\{f_{1}(\sigma)\}\bigcup_{\alpha\in A}\{g_{1}(\alpha)\}$, access $f_{i}(\sigma)$ for $i=2,...,k$ and compare the resulting vector value to $g(\alpha)$ for all $\alpha\in A$. Since $|A|\leq\dim\sigma+1\leq d+1$, where $d:=\dim K$, each of these operations is done in $O(dk)$ time at most.
@@ 5. $\mathrm{(5)}$
    The cost of add_cofacets when applied to $\sigma$ depends on its number of cofacets, which may be quite large when $K$ is an arbitrary simplicial complex. Nonetheless, we know that each $\beta\in L_{u}$ will be checked by add_cofacets at most as many times as the number of facets of $\beta$, which is bounded by $d+1$. Hence, we can approximate of the cost of add_cofacets is $O(d)$ for each simplex in $L_{u}$.
@@
$O\left(|f(K)|\log\lambda^{\lambda}+dk|K|\right)$
Finally, by adding the cost of LevelSets to that of ExpandMDM, we obtain a total cost for GenerateMDM of $O\left(|K|+|f(K)|\log|f(K)|+|f(K)|\log\lambda^{\lambda}+dk|K|\right)$, which can be rewritten as
@@ 3. $\mathrm{(3)}$
Second, being able to process a map $f$ that is not injective means that the algorithm may also be used to generate an arbitrary mdm function and its gradient field on a given complex. Indeed, if we consider $f$ to be a constant map, such as the one in Figure 5a, then there exists a unique level set $L_{u}$ in $K$, which is the whole complex. Thus, GenerateMDM calls ExpandMDM only once and it processes the whole complex $K=L_{u}$ at one time. Also, we see that in this case, $|f(K)|=1$ and $\lambda:=\max_{u\in f(K)}|L_{u}|=|K|$. Hence, from Section 4.2, we have that GenerateMDM takes at most $O(dk|K|+|K|\log|K|)$ time, or simply $O(|K|\log|K|)$ time if we assume $d$ and $k$ to be small.
@@
Experimentally, we see that a mdm function generated this way is often optimal, in the sense that its number of critical simplices of index $p$ equals the $p^{\text{th}}$ Betti number of $K$ (see Section 7.1). However, for large datasets, GenerateMDM does not always minimize the number of critical simplices of the output, which is to be expected since computing an optimal discrete gradient field on general simplicial complexes is $\mathcal{NP}$-hard [38, 49]. Still, the algorithms designed to find an optimal gradient field require either $\dim K\leq 2$ [49] or $K$ to have a manifold structure [10], while heuristics which aims to do this for general simplicial complexes run in $O(|K|^{2})$ time [50]. This suggests that in addition to being well adapted to the context of multipersistent homology, our approach is efficient for generating an arbitrary discrete gradient field with an almost minimal number of critical simplices on a general simplicial complex.

### v3

@@ 3.3. Critical components
Let $R$ be a relation on the set ${\mathcal{C}}$ of critical simplices of a mdm function and consider its transitive closure $\bar{R}$. Namely, we have $\sigma\bar{R}\tau$ if there is a sequence $\sigma=\sigma_{0},\sigma_{1},...,\sigma_{n}=\tau\in{\mathcal{C}}$ such that $\sigma_{i-1}R\sigma_{i}$ for each $i=1,...,n$. If $R$ is reflexive and symmetric, by definition, it follows that $\bar{R}$ is an equivalence relation, so it can be used to partition ${\mathcal{C}}$. We then call a critical component (with respect to $\bar{R}$) an equivalence class of $\bar{R}$. As an example, if $R$ is such that $\sigma R\tau\Leftrightarrow\sigma\leq\tau\text{ or }\sigma\geq\tau$, then $\bar{R}$ yields the partition of ${\mathcal{C}}$ into connected components (with respect to the Alexandrov topology).
As mentioned above, connectedness is too strict of a criteria to define critical components. Thus, we consider the following definition, first introduced as Definition 7.3 in [14]. It makes use of the idea of dynamical connectedness between critical simplices. More precisely, for two critical simplices $\sigma$ and $\tau$, we write $\sigma\rightarrowtriangle_{g}\tau$ when $\sigma$ is connected to $\tau$ in the flow induced by the gradient field ${\mathcal{V}}$ of $g$. This is the case, for example, when $\sigma\geq\tau$ or when there is a ${\mathcal{V}}$-path going from a face of $\sigma$ to a coface of $\tau$. For a formal definition of the flow of ${\mathcal{V}}$, see [14].
Definition 3.5.
Let $g:K\rightarrow{\mathbb{R}}^{k}$ be a mdm function and ${\mathcal{C}}$ be the set of its critical simplices. Consider the relation $R_{g}$ defined on ${\mathcal{C}}$ such that $\sigma R_{g}\tau$ if
@@ 2. $\mathrm{(2)}$
Then, $\sim_{g}:=\bar{R}_{g}$ is an equivalence relation, and we define a critical component of $g$ (with respect to $\sim_{g}$) as an equivalence class of $\sim_{g}$.
@@
For toy examples, defining critical components using $\sim_{g}$ yields desirable results, as argued in [14]. For instance, the mdm function from Example 3.4 has two critical components with respect to $\sim_{g}$, represented in red and orange in Figure 3. In Section 6.3, we will see how this definition of critical components of a mdm function performs on a larger scale. Also, a new definition of critical components will be proposed, Definition 6.1, which is adapted specifically to mdm functions which are compatible with a given multifiltering function.
@@ 4.1. Description
First of all, apart from the complex $K$ and the map $f$, the main algorithm takes two input parameters: an injective real-valued map $I:K\rightarrow{\mathbb{R}}$, called an indexing map, and a small parameter $\epsilon>0$. The indexing map $I$ is used to order the simplices of $K$ while the parameter $\epsilon$ is used to ensure the mdm function $g$ computed is such that, for all $\sigma\in K$ and all $i=1,...,k$,
@@
$w_{1}$ $=\max\left(\{f_{1}(\sigma)\}\cup\{g_{1}(\alpha)\ |\ \alpha\in A\}\right),$
@@
7:     $w_{1}=\max\left(\{f_{1}(\sigma)\}\cup\{g_{1}(\alpha)\ |\ \alpha\in A\}\right)$
@@
Remark 4.1.
The procedure ComputeG returns a value using all facets $\alpha$ of a simplex $\sigma$, where each $\alpha$ may or may not be in the same level set as $\sigma$. That being said, ExpandMDM only pairs together simplices that belong in a same level set. Hence, if we were to adapt the main algorithm GenerateMDM in order to produce only a gradient field, it would be fairly easy to parallelize the procedure by calling ExpandMDM separately on each level set.
4.2. Correctness
We now explain broadly why the proposed algorithms produce the desired result. The formal proofs are available in Appendix A.
First, we consider the routine ExpandMDM, described in Algorithm 2. When ExpandMDM is called, we see that all simplices with 00 or $1$ facet in $L_{u}$, facet which is necessarily unprocessed at this point, immediately enter PQzero or PQone. These simplices eventually get processed, so the number of unprocessed facets of their cofacets gradually decreases. When the number of unprocessed facets of a simplex reaches $1$, it is added to PQone by the subroutine add_cofacets, so it can also eventually be processed. In other words, ExpandMDM first processes the lower-dimensional simplices and slowly works its way up to higher-dimensional simplices until all of $L_{u}$ is processed. Since the processed simplices never re-enter PQzero or PQone, we obtain the next result.
{re-prop}
[] For all $u\in f(K)$, each simplex in $L_{u}$ is processed exactly once by Algorithm 2.
Hence, since Proposition 4.2 stands for every level set $L_{u}\subset K$ and each $L_{u}$ is processed exactly once by the main routine GenerateMDM, presented in Algorithm 1, we deduce that every $\sigma\in K$ is processed exactly once by GenerateMDM.
Moreover, recall that if two simplices $\alpha<\sigma$ are part of a same priority queue, $\alpha$ has priority over $\sigma$ because $I(\alpha)<I(\sigma)$. Hence, from the reasoning leading to Proposition 4.2, we deduce the following key observation: the facets of a given simplex $\sigma$ are always processed either before $\sigma$, or at the same time by being paired by ExpandMDM. Therefore, every time ComputeG is called in Algorithm 2 to define some value $g(\sigma)$, and optionally $g(\tau)$ when $\sigma$ is paired with a facet $\tau$, we know the value $g(\alpha)$ has been defined beforehand for all facets in $A=\{\text{facets }\alpha<\sigma\ |\ \alpha\neq\tau\}$. Thus, ComputeG always returns a value $w\in{\mathbb{R}}^{k}$ without error. More specifically, we can characterize $g$ as follows.
{re-prop}
[] Consider function ComputeG described in Algorithm 3.
  1. $\mathrm{(1)}$
    When called in Algorithm 2, ComputeG returns a value $w\in{\mathbb{R}}^{k}$ without error, so the output dictionary $g$ of Algorithm 1 is such that $g(\sigma)$ is well defined for all $\sigma\in K$.
  2. $\mathrm{(2)}$
  3. $\mathrm{(3)}$
    For all $\sigma\in K$, we have
    $g_{1}(\sigma)$ $=f_{1}(\sigma)+m\delta\text{ for some }m\in{\mathbb{N}},$
    $g_{i}(\sigma)$ $=f_{i}(\sigma)\text{ for each }i=2,...,k$
Now, we could show that the dictionary ${\mathcal{V}}$ output by GenerateMDM pairs together two simplices $\tau<\sigma$ if and only if $\sigma$ and $\tau$ belong in a same level set $L_{u}$ and $g(\tau)=g(\sigma)$. Otherwise, when $\tau<\sigma$ are not paired, we have $g(\tau)\precneqq g(\sigma)$. Since every simplex is processed exactly once, it belongs in at most one pair defined by ${\mathcal{V}}$, and we can deduce that $g$ is in fact a mdm function. Also, from part $\mathrm{(3)}$ of Proposition 4.2, we see that
because $\delta$ is defined in Algorithm 1 so that $\delta\leq\frac{\epsilon}{|K|}$. Thus, we conclude that GenerateMDM produces the desired result, as stated in the following theorem.
{re-thm}
[] For any simplicial complex $K$, multifiltering function $f$, admissible indexing map $I:K\rightarrow{\mathbb{R}}$ and $\epsilon>0$ given as input in GenerateMDM, the outputs $g$ and ${\mathcal{V}}$ correspond to a $f$-compatible mdm function $g:K\rightarrow{\mathbb{R}}^{k}$ and its gradient field ${\mathcal{V}}:K\nrightarrow K$ such that, for all $\sigma\in K$,
$\lVert g(\sigma)-f(\sigma)\rVert<\epsilon$
[…]
Furthermore, we can see GenerateMDM as an algorithm which outputs a mdm function $g$ from an input admissible map $f$ by dividing its level sets into pairs and singletons. When $f$ is itself mdm, the connected components of its level sets are all of cardinality $1$ or $2$, so GenerateMDM simply returns $g:=f$.
{re-prop}
[] Let $g:K\rightarrow{\mathbb{R}}^{k}$ be the mdm function produced by GenerateMDM when given $f:K\rightarrow{\mathbb{R}}^{k}$ as input. The function $f$ is itself mdm if and only if $f=g$.
@@ 4.3. Complexity analysis
  • The index mapping $I$ is precomputed. As explained previously at the beginning of Section 4.1, the computation of $I$ can be included in the implementation of $K$.
@@
Moreover, the complexity of GenerateMDM depends greatly on the execution of ExpandMDM. For each level set $L_{u}$, we know from Section 4.2 that each $\sigma\in L_{u}$ is added to each priority queue PQzero and PQone at most once and is processed exactly once. We compute the cost of processing each $\sigma\in L_{u}$ by considering each step of its process separately.
@@ 3. $\mathrm{(3)}$
    From Algorithm 3, we see that to define $g(\sigma)$, in the worst case, ComputeG needs to compute $\max\left(\{f_{1}(\sigma)\}\cup\{g_{1}(\alpha)\ |\ \alpha\in A\}\right)$, access $f_{i}(\sigma)$ for $i=2,...,k$ and compare the resulting vector value to $g(\alpha)$ for all $\alpha\in A$. Because $|A|\leq\dim\sigma+1\leq d+1$, where $d:=\dim K$, each of these operations is done in $O(dk)$ time at most.
@@ 5. $\mathrm{(5)}$
    The cost of add_cofacets when applied to $\sigma$ depends on its number of cofacets, which may be quite large when $K$ is an arbitrary simplicial complex. Nonetheless, we know that each $\gamma\in L_{u}$ will be checked by add_cofacets at most as many times as the number of facets of $\gamma$, which is bounded by $d+1$. Hence, we can approximate of the cost of add_cofacets is $O(d)$ for each simplex in $L_{u}$.
@@
$O\left(|f(K)|\log\lambda^{\lambda}+dk|K|\right).$
Finally, by adding the costs of LevelSets and ExpandMDM, we obtain a total cost for GenerateMDM of $O\left(|K|+|f(K)|\log|f(K)|+|f(K)|\log\lambda^{\lambda}+dk|K|\right)$, which can be rewritten as
@@ 3. $\mathrm{(3)}$
Second, being able to process a map $f$ that is not injective means that the algorithm may also be used to generate an arbitrary mdm function and its gradient field on a given complex. Indeed, if we consider $f$ to be a constant map, such as the one in Figure 5a, then there exists a unique level set $L_{u}$ in $K$, which is the whole complex. Thus, GenerateMDM calls ExpandMDM only once and it processes the whole complex $K=L_{u}$ at one time. Also, we see that in this specific case, $|f(K)|=1$ and $\lambda:=\max_{u\in f(K)}|L_{u}|=|K|$. Hence, from Proposition 4.2 in Section 4.3, we have that GenerateMDM takes at most $O(dk|K|+|K|\log|K|)$ time, or simply $O(|K|\log|K|)$ time if we assume $d$ and $k$ to be small.
@@
Experimentally, we see that a mdm function generated this way is often perfect, meaning that its number of critical simplices of index $p$ equals the $p^{\text{th}}$ Betti number of $K$ (see Section 6.1). However, for large datasets, GenerateMDM does not always minimize the number of critical simplices of the output, which is to be expected since computing an optimal discrete gradient field on general simplicial complexes is $\mathcal{NP}$-hard [39, 50]. Still, the algorithms designed to find an optimal gradient field require either $\dim K\leq 2$ [50] or $K$ to have a manifold structure [10], while heuristics which aims to do this for general simplicial complexes run in $O(|K|^{2})$ time [51]. This suggests that in addition to being well adapted to the context of multipersistent homology, our approach is efficient for generating an arbitrary discrete gradient field with an almost minimal number of critical simplices on a general simplicial complex.